### v2

# DOFA-CLIP: Multimodal Vision–Language Foundation Models for Earth Observation

Abstract
Earth observation (EO) spans a broad spectrum of modalities, including optical, radar, multispectral, and hyperspectral data, each capturing distinct environmental signals. However, current vision-language models in EO, particularly CLIP-based variants, remain confined to individual modalities, limiting generalization and scalability across diverse tasks. We present DOFA-CLIP (Dynamic-One-For-All CLIP), a unified vision-language foundation model that dynamically adapts to EO modalities with flexible spectral configurations through a single Transformer backbone. Our approach introduces three key contributions: 1) the construction of GeoLangBind-2M, a large-scale EO image–text dataset covering six heterogeneous modalities with rich natural language descriptions; 2) a novel training strategy called VECT (Vision-models Enhanced Contrastive Text-image pretraining), which enhances the spatial awareness of CLIP features with multiple vision foundation models; and 3) a Modality-aware Knowledge Agglomeration (MaKA) module that refines feature distillation with modality-specific awareness. DOFA-CLIP achieves state-of-the-art zero-shot performance across a wide range of EO benchmarks, including unseen modalities and a diverse number of input spectral bands. Together, these contributions establish a scalable foundation for multimodal EO understanding and open new avenues for integrating heterogeneous EO data with large language models. Code and datasets will be released.
Introduction
Figure 1: DOFA-CLIP employs a single unified vision Transformer to align heterogeneous EO data modalities using language as the grounding signal for cross-modal understanding and zero-shot generalization.
Earth observation (EO) technologies have undergone remarkable advancement in recent years, generating vast amounts of multimodal geospatial data spanning different sensors and scales (Xia et al. 2018; Zhu et al. 2017; Xiong et al. 2024c). In response to the exponential growth of remote sensing data, researchers have proposed foundation models (Cong et al. 2022; Reed et al. 2023; Wang et al. 2024a, 2023a; Stewart et al. 2023; Hong et al. 2024) pretrained on large-scale EO data for multiple downstream tasks. However, these models typically focus on individual EO data modality, limiting their ability to leverage complementary characteristics across different sensor types (Zhu et al. 2024).
To address this limitation, recent works such as OFA-Net (Xiong et al. 2024b), DOFA (Xiong et al. 2024a), AnySat (Astruc et al. 2024a), OmniSat (Astruc et al. 2024b), and FoMo-Net (Bountos, Ouaknine, and Rolnick 2023) have introduced foundation models capable of learning from multiple EO modalities within a unified framework. These approaches have demonstrated promising results in multimodal EO representation learning. However, their focus remains on unsupervised visual pretraining, lacking a structured way to align EO data with language for better interpretability and retrieval. Meanwhile, multimodal large language models like GPT (OpenAI 2023), LLaVA (Liu et al. 2023), and Pixtral (Agrawal et al. 2024) have shown remarkable capabilities in handling diverse and complex tasks. Vision-language foundation models, such as CLIP (Radford et al. 2021) and SigLIP (Zhai et al. 2023), have played a crucial role as vision encoders of such multimodal large language models. Inspired by these advances, EO-specific adaptations like RemoteCLIP (Liu et al. 2024), RS5M (Zhang et al. 2024b), RS-CLIP (Li et al. 2023), and SkyScript (Wang et al. 2024c) have attempted to bring vision-language contrastive learning to remote sensing. However, existing CLIP models for EO have limitations: 1) These models are usually limited to high-resolution RGB imagery and cannot handle modalities like multi/hyper-spectral or SAR, which are vital for many applications. 2) Compared with vision models like DINOv2 (Oquab et al. 2023), CLIP models are unsuitable for dense understanding tasks due to the lack of spatial awareness.
Training a single powerful vision–language model capable of handling heterogeneous EO data modalities raises several core research questions (RQ). These include:
[…]
    RQ1: How can we build a large-scale, semantically rich image–text dataset for diverse heterogeneous EO modalities?
[…]
    RQ2: How can we efficiently train large CLIP models for EO with better spatial awareness?
[…]
    RQ3: How can a single CLIP model handle heterogeneous EO data while preserving modality-specific characteristics?
To address such challenges, we propose Dynamic-One-For-All CLIP (DOFA-CLIP), a unified multimodal vision–language foundation model that aligns diverse EO modalities to the language space using a single Transformer, as shown in Fig. 1. Our contributions are as follows:
  1. 1.
    We introduce GeoLangBind-2M, a large-scale EO dataset comprising two million image–text pairs across six heterogeneous modalities, providing high-quality text descriptions for multimodal vision–language training.
  2. 2.
    We propose Vision-models Enhanced Contrastive Text-image pretraining (VECT), a simple yet effective training strategy that enhances the fine-grained spatial awareness of CLIP models with open vision foundation models.
  3. 3.
    We develop the Modality-aware Knowledge Agglomeration (MaKA) module to support modality-adaptive representation distillation within a unified Transformer.
  4. 4.
    We conduct extensive evaluations across zero-shot classification and cross-modal retrieval tasks, demonstrating state-of-the-art performance and better spatial awareness on both seen and unseen EO modalities.
Figure 2: Visualization of data samples from GeoLangBind-2M. The dataset includes imagery from six different sensors and modalities: Sentinel-2 multispectral, Sentinel-1 SAR, EnMAP hyperspectral, elevation maps, infrared imagery, and aerial imagery. Each sample is paired with textual descriptions capturing key land cover types, objects, and geographic features.
Related work
Vision foundation models in Earth observation
Existing pretrained models predominantly focus on RGB data, such as GFM (Mendieta et al. 2023), Scale-MAE (Reed et al. 2023), and Cross-Scale MAE (Tang et al. 2024). Others specialize in multispectral imagery, with FG-MAE (Wang et al. 2023b) and SatMAE (Cong et al. 2022) tailored for Sentinel-2 data, while SSL4EO-L (Stewart et al. 2023) is trained on Landsat imagery. SpectralGPT (Hong et al. 2024) leverages a 3D generative transformer for spectral data analysis. CROMA (Fuller, Millard, and Green 2023) and DeCUR (Wang et al. 2024b) use self-supervised multimodal learning to learn features across modalities. Satlas (Bastani et al. 2023) assembles a large-scale multi-sensor dataset, pretraining separate models for each sensor. DOFA (Xiong et al. 2024a) introduces a neural plasticity-inspired hypernetwork that adapts to different sensor wavelengths, enabling joint training on five distinct sensors. OmniSat (Astruc et al. 2024b) proposes to fuse features across multiple EO modalities to learn robust multimodal representations without labeled data. SkySense (Guo et al. 2024) utilizes separate visual encoders for different data modalities and designs a dedicated module for multimodal fusion.
Vision-language foundation models
Zero-shot foundation models like CLIP (Radford et al. 2021) have significantly advanced computer vision. Building upon CLIP, SigLIP (Zhai et al. 2023) introduces a pairwise sigmoid loss function to enlarge the training batch and enhance the model performance. Zero-shot foundation models have also been investigated in EO applications. SkyScript (Wang et al. 2024c) constructs a large-scale image–text dataset by linking remote sensing images with OpenStreetMap (OpenStreetMap contributors 2017) semantics. RemoteCLIP (Liu et al. 2024) introduces an image–text dataset by generating captions from existing annotated datasets. GeoRSCLIP (Zhang et al. 2024b) introduces RS5M, an image–text dataset with 5M pairs, created by filtering existing datasets and generating captions for label-only remote sensing data. Despite these advances, existing vision–language models mainly focus on RGB data and have limitations in learning representations from multiple EO data modalities. To unify multiple data modalities, ImageBind (Girdhar et al. 2023) and TaxaBind (Sastry et al. 2025) align multiple modalities within a shared embedding space. LanguageBind (Zhu et al. 2023) uses language as a universal intermediary for modality alignment. In the era of Large Language Models (LLMs), methods such as GeoChat (Kuckreja et al. 2024), SkyEyeGPT (Zhan, Xiong, and Yuan 2025), and TeoChat (Irvin et al. 2024) integrate CLIP-based vision models with LLMs to unify diverse tasks and achieve impressive performance.
GeoLangBind-2M Dataset
Table 1: Summary of dataset sources of GeoLangBind-2M. * indicates image–text datasets created in this work. SAR = synthetic aperture radar, MSI = multispectral imagery, HSI = hyperspectral imagery, and IR = infrared.
Image–Text Datasets Knowledge Domain # Samples
Seg4 General segmentation 41,172
Det10 General detection 110,800
Flair2-RGB-caption* Land cover analysis 61,711
VRS-train General detection 20,264
SkyScript dataset OpenStreetMap tags 1,518,888
SAR Datasets
MMflood-caption* Flood mapping 6,181
SAR-ship-caption* Ship detection 5,984
ChatEarthNet-SAR-caption* Land cover analysis 95,620
MSI/HSI Datasets
ChatEarthNet-S2-caption* Land cover analysis 95,620
NLCD-hyper-caption* Land cover analysis 15,000
Elevation Dataset
Flair2-Elevation-caption* Digital Elevation Model 61,711
Infrared Dataset
IR-ship-caption* Ships in infrared imagery 18,032
Total 2,050,983
In response to RQ1, we propose GeoLangBind-2M, a large-scale image-text dataset covering a broad range of modalities including RGB, SAR, multispectral, hyperspectral, infrared, and elevation data, as listed in Table 1. Each image is paired with a high-quality text annotation, as illustrated in Fig. 2. All datasets are implemented using the TorchGeo library (Stewart et al. 2022). In the following, we introduce the composition of the dataset.
RGB image–text datasets
High-resolution RGB images with textual descriptions form a significant portion of GeoLangBind-2M. We incorporate parts of the RemoteCLIP (Liu et al. 2024) dataset: Seg4 and Det10. We also include the training split of VRSBench (Li, Ding, and Elhoseiny 2024), which aggregates multiple object detection datasets. SkyScript (Wang et al. 2024c) is a global image–text dataset for EO, from which we include the top 30% highest quality samples. In addition to existing datasets, we construct the Flair2-RGB-caption (Garioud et al. 2023) dataset by leveraging the Pixtral-12B (Agrawal et al. 2024) model, using semantic information as prompts to generate captions for land cover features.
SAR image–text datasets
To enhance the dataset with SAR–text pairs, we create MMFlood-SAR-caption (Montello, Arnaudo, and Rossi 2022), which contains captions emphasizing water extent and flooded regions (more details in supplementary material). Additionally, the SAR-ship-caption dataset includes captions describing ships in SAR images. ChatEarthNet-SAR-caption further extends the collection by covering diverse SAR scenes from Sentinel-1, focusing on land cover types.
Multispectral, hyperspectral, and elevation image–text datasets
We construct ChatEarthNet-S2-caption by summarizing long, detailed descriptions from the ChatEarthNet dataset (Yuan et al. 2024). This dataset contains Sentinel-2 multispectral imagery covering more than 10 land cover types. Furthermore, we introduce NLCD-hyper-caption (Braham et al. 2024), a hyperspectral dataset with 200+ EnMAP bands (Guanter et al. 2015). Like Flair2-RGB-caption, we use Pixtral-12B to generate rich captions, leveraging annotated land cover maps for enhanced descriptiveness. Similarly, we also generate captions using Pixtral-12B to construct the Flair2-Elevation-caption dataset. Further details on the construction of GeoLangBind-2M are provided in the Supplementary materials.
Figure 3: DOFA-CLIP aims to align diverse Earth observation data modalities into the language feature space. Vision-models Enhanced Contrastive Text-image pretraining (VECT) is proposed to enhance fine-grained image understanding. Modality-aware knowledge agglomeration (MaKA) is designed to transfer knowledge from heterogeneous sensors. The wavelength-aware dynamic patch embedding module processes multimodal images dynamically.
Dynamic-One-For-All CLIP
Wavelength-aware dynamic encoder
To handle the diversity of spectral bands across modalities, we adopt the dynamic encoder architecture from DOFA (Xiong et al. 2024a), as illustrated in Fig. 3. Given an input image $\mathbf{X}\in\mathbb{R}^{C\times H\times W}$ with $C$ spectral channels, we define its corresponding central wavelengths as $\mathbf{\lambda}\in\mathbb{R}^{C}$. For modalities without well-defined wavelengths, such as elevation, we assign default wavelengths equivalent to those of the RGB channels. First, a 1D sine-cosine positional encoding is applied to embed these wavelengths into a higher-dimensional space: $\mathbf{V}_{\lambda}=\text{PE}(\mathbf{\lambda})\in\mathbb{R}^{C\times d_{\lambda}},$ where $d_{\lambda}$ is the embedding dimension. The encoded wavelengths $\mathbf{V}_{\lambda}$ are then transformed using two fully-connected layers with residual connections. Next, a lightweight Transformer (Vaswani et al. 2017) encoder processes the transformed wavelength embeddings. Specifically, we concatenate the transformed embeddings $\mathbf{V^{\prime}}_{\lambda}$ with the learnable weight query tokens $\mathbf{Q}_{w}$ and a learnable bias query token $\mathbf{Q}_{b}$ as input to the Transformer encoder. After the Transformer layer, the output features are passed through two fully-connected layers to generate the dynamic convolutional weights $\mathbf{M}_{w}$ and biases $\mathbf{M}_{b}$. Finally, the dynamically generated weights and biases are used for patch embedding to handle EO data modalities with varying spectral channels.
Vision-models enhanced CLIP
To address RQ2, we propose VECT to enhance the spatial coherence of CLIP features by knowledge distillation from open vision models. Specifically, in addition to the contrastive loss, VECT leverages a feature-matching loss to align the intermediate features of DOFA-CLIP with features from multiple pre-trained vision models: SigLIP (Zhai et al. 2023), DINOv2 (Oquab et al. 2023), and ViT (Dosovitskiy et al. 2020). As illustrated in Fig. 3, we extract three RGB channels from multispectral or hyperspectral images as input to the teacher model. The student features $\mathbf{F}_{s}\in\mathbb{R}^{d_{t}\times H^{\prime}\times W^{\prime}}$ are then matched to the teacher’s features $\mathbf{F}_{t}\in\mathbb{R}^{d_{t}\times H^{\prime}\times W^{\prime}}$ using a combination of $L_{1}$ loss, cosine embedding loss, and mean squared error (MSE) loss:
$L_{\text{match}}=L_{L_{1}}(\mathbf{F}_{s},\mathbf{F}_{t})+L_{\text{mse}}(\mathbf{F}_{s},\mathbf{F}_{t})+L_{\text{cos}}(\mathbf{F}_{s},\mathbf{F}_{t}).$ (1)
DOFA-CLIP is trained on GeoLangBind-2M using both the pairwise sigmoid loss $L_{\text{siglip}}$ for contrastive learning and distillation loss for feature alignment. The total loss function is:
$L=L_{\text{siglip}}+\alpha_{s}L_{\text{match}}^{\text{siglip}}+\alpha_{d}L_{\text{match}}^{\text{dinov2}}+\alpha_{v}L_{\text{match}}^{\text{vit}},$ (2)
where $\alpha_{s},\alpha_{d},$ and $\alpha_{v}$ are balancing factors for feature-matching losses of SigLIP, DINOv2, and ViT, respectively.
Modality-aware knowledge agglomeration
Considering RQ3, we introduce the modality-aware knowledge agglomeration (MaKA) module, which is used to translate DOFA-CLIP features in a modality-aware manner to facilitate effective feature matching in VECT. MaKA comprises three key components: a wavelength-aware prompt generator, a modality-aware conditional layer normalization, and a feature projector.
1. Wavelength-aware prompt generator
Each modality is associated with a set of wavelengths, which we encode using a sine-cosine positional encoding function. Given input wavelengths $\mathbf{\lambda}\in\mathbb{R}^{C}$, we first obtain their positional embeddings $\mathbf{V}_{\lambda}\in\mathbb{R}^{C\times d_{\lambda}}$. To generate the modality-aware prompt vector, we apply a linear projection $\mathbf{W}_{\text{proj}}\in\mathbb{R}^{d\times d_{\lambda}}$ to each row of $\mathbf{V}_{\lambda}$, then compute the mean over the $C$ wavelengths:
$\mathbf{V}_{p}=\frac{1}{C}\sum_{i=1}^{C}\bigl{(}\mathbf{W}_{\text{proj}}\,\mathbf{V}_{\lambda_{i}}\bigr{)},$ (3)
where $\mathbf{V}_{p}\in\mathbb{R}^{d}$ is the fused modality embedding. This final embedding vector $\mathbf{V}_{p}$ captures modality-specific information derived from the provided wavelengths.
2. Modality-aware layer normalization
Given the feature map from the last layer of the DOFA-CLIP model, we first interpolate it to match the teacher’s feature resolution, denoted as $\mathbf{F}\in\mathbb{R}^{d\times H^{\prime}\times W^{\prime}}$, omitting the batch dimension for simplicity. With $\mathbf{F}$ and a modality prompt $\mathbf{V}_{p}\in\mathbb{R}^{d}$, our goal is to perform layer normalization that dynamically adapts to the modality. To achieve this, we introduce modality-aware conditional layer normalization as follows.
First, we derive two learnable modulation vectors, $\boldsymbol{\gamma},\boldsymbol{\beta}\in\mathbb{R}^{d}$, from the modality prompt via a linear transformation:
$\begin{bmatrix}\boldsymbol{\gamma}\\[6.0pt] \boldsymbol{\beta}\end{bmatrix}\;=\;\mathbf{W}_{\text{prompt}}\,\mathbf{V}_{p}\;\in\;\mathbb{R}^{2d},$ (4)
where $\mathbf{W}_{\text{prompt}}\in\mathbb{R}^{2d\times d}$. We then split the output into $\boldsymbol{\gamma},\boldsymbol{\beta}\in\mathbb{R}^{d}$ and reshape them to broadcast over the spatial dimensions:
[…]
Let $\mathrm{LN}(\mathbf{F})$ denote the standard channel-wise layer normalization of $\mathbf{F}$. The modality-aware layer normalization is then defined as:
$\mathbf{F}^{\prime}\;=\;\mathrm{LN}(\mathbf{F})+\boldsymbol{\gamma}\odot\mathrm{LN}(\mathbf{F})\;+\;\boldsymbol{\beta},$ (6)
where $\odot$ denotes the Hadamard product. Note that we use the modality-aware layer normalization to learn the residual information specific to each modality.
3. Projector module
A convolutional layer is applied to adjust the feature representation $\mathbf{F}^{\prime}$ to channel dimensions $d_{t}$ of the corresponding teachers. The final representation is then used to compute the feature-matching loss for model distillation. This design ensures that the learned features are both aligned and modality-specific, enhancing robust adaptation across different input modalities.
To address the data imbalance issue, we employ a weight space merging strategy (Ilharco et al. 2022). Since RGB data is the most dominant modality, we split the dataset into two parts: RGB data and non-RGB data. We first train two separate DOFA-CLIP models on these subsets. Then, we merge their weights to obtain a unified model that can handle both RGB and other modalities. Since the original SigLIP model is trained on massive data, we also merge the original SigLIP weights with our model. For clarity, let $\theta_{\text{siglip}}$ be the original SigLIP model weights, $\theta_{\text{rgb}}$ be the DOFA-CLIP model weights trained on the RGB subset, $\theta_{\text{others}}$ be the DOFA-CLIP model weights trained on the other five modalities. In the first step, we merge $\theta_{\text{siglip}}$ and $\theta_{\text{rgb}}$ using a simple linear weight merging strategy:
$\theta^{*}=(1-m_{1})\,\theta_{\text{siglip}}+m_{1}\,\theta_{\text{rgb}},$ (7)
where $m_{1}$ controls the weighting ratio between the two models. Next, we merge the intermediate weights $\theta^{*}$ with $\theta_{\text{others}}$ to obtain the final model:
$\theta=(1-m_{2})\,\theta^{*}+m_{2}\,\theta_{\text{others}},$ (8)
where $m_{2}$ determines the contribution of the non-RGB modalities. We conduct ablation studies by varying $m_{1}$ and $m_{2}$ to search for optimal weighting ratios. While more advanced merging methods exist (Akiba et al. 2025), we use a linear approach to assess its effectiveness on heterogeneous EO data, leaving sophisticated strategies for future work.
Table 2: Zero-shot comparison of various models on scene and fine-grained classification tasks. Bold values indicate the best performance.
ViT Model Scene classification Fine-grained classification
SkyScript AID EuroSAT fMoW Million-AID PatternNet RESISC RSI-CB Avg. Roof shape Smoothness Surface
Base CLIP-original 40.16 69.55 32.11 17.62 57.27 64.09 65.71 41.26 49.66 31.50 26.80 61.36
Human-curated captions 40.03 71.05 33.85 18.02 57.48 66.56 66.04 42.73 50.82 28.50 27.80 60.91
RemoteCLIP 27.06 87.05 30.74 11.13 46.26 56.05 67.88 44.55 49.09 30.50 21.00 43.86
CLIP-laion-RS 40.77 69.55 37.63 19.16 56.59 64.79 64.63 41.79 50.59 28.83 27.60 62.27
SkyCLIP-50 52.98 70.90 33.30 19.24 62.69 72.18 66.67 46.20 53.02 26.00 38.00 67.73
DOFA-CLIP-B-224 70.39 77.60 52.30 20.17 64.91 76.68 67.21 49.53 59.85 44.33 20.00 72.73
Large CLIP-original 55.06 69.25 41.89 26.19 57.88 71.39 66.70 43.02 53.76 37.50 25.40 42.73
Human-curated captions 56.09 72.95 41.96 26.33 58.47 74.86 68.70 44.60 55.41 37.00 26.60 40.00
RemoteCLIP 34.40 70.85 27.81 16.77 47.20 61.91 74.31 50.79 49.99 34.33 34.20 55.45
CLIP-laion-RS 58.81 71.70 54.30 27.21 60.77 72.68 71.21 48.21 57.82 40.50 37.60 53.41
SkyCLIP-20 67.94 71.95 53.63 28.04 65.68 78.62 70.70 50.03 59.98 44.83 26.80 61.36
SkyCLIP-30 69.08 72.15 52.44 27.77 66.40 79.67 70.77 50.19 59.99 46.17 30.80 64.32
SkyCLIP-50 70.89 71.70 51.33 27.12 67.45 80.88 70.94 50.09 59.93 46.83 35.80 67.50
DOFA-CLIP-L-384 76.83 75.50 59.04 29.10 70.16 80.17 73.15 51.62 64.45 61.83 26.00 81.36
Table 3: Zero-shot comparison on GeoBench datasets with non-RGB data. Bold values indicate the best performance.
Model (Base version) m-bigearthnet m-so2sat m-forestnet
Precision Recall F1 RGB Sentinel-2 (5 bands) RGB Landsat-8 (5 bands)
SigLIP 45.34 12.36 16.82 12.88 - 8.16 -
RemoteCLIP 33.82 20.35 18.84 10.75 - 8.46 -
SkyCLIP-50 39.91 19.58 20.32 11.97 - 10.78 -
DOFA-CLIP-B (224) 47.70 20.37 23.69 17.95 14.60 13.60 17.02
Experiments
Implementation details
DOFA-CLIP is trained on 8 NVIDIA A100 40GB GPUs for the base model and 8 NVIDIA A100 80GB GPUs for the large version. For the base version, we use a ViT-Base architecture (patch size 16) pretrained on WebLI (Chen et al. 2022). The teacher models are ViT-Base-SigLIP (224), DINOv2-ViT-L/14, and ViT-Huge (ImageNet). The large version is based on SoViT-400m (Alabdulmohsin et al. 2023) (patch size 14, input size 384), with teacher models ViT-So400m-SigLIP (384), DINOv2-ViT-L/14, and ViT-Large (ImageNet). Training is conducted for 20 epochs using AdamW (learning rate 5e-4, weight decay 1e-7).
Performance comparison
We follow the evaluation protocols of Wang et al. (2024c) and evaluate the models via zero-shot classification tasks. Specifically, the following eleven image scene classification datasets are used: AID (Xia et al. 2017), EuroSAT (Helber et al. 2019), fMoW (Christie et al. 2018), Million-AID (Long et al. 2021), PatternNet (Zhou et al. 2018), NWPU-RESISC45 (Cheng, Han, and Lu 2017), RSI-CB256 (Li et al. 2017), as well as the fine-grained attribute classification datasets. In addition, we further compare these models on some datasets from the GEO-Bench suite (Lacoste et al. 2024): m-bigearthnet (Sumbul et al. 2019), m-so2sat (Zhu et al. 2019), and m-forestnet (Irvin et al. 2020).
Table 4: Recall@1, recall@5, and recall@10 for cross-modal retrieval on three benchmark datasets. Bold values indicate the best performance.
RSICD RSITMD UCM-caption
Image to Text Text to Image Image to Text Text to Image Image to Text Text to Image
Model R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10 R@1 R@5 R@10
RSICD, RSITMD, and/or UCM-caption seen in training
AMFMN (Yuan et al. 2022a) 5.39 15.08 23.40 4.90 18.28 31.44 10.63 24.78 41.81 11.51 34.69 54.87 16.67 45.71 68.57 12.86 53.24 79.43
LW-MCR-u (Yuan et al. 2021) 4.39 13.35 20.29 4.30 18.85 32.34 9.73 26.77 37.61 9.25 34.07 54.03 18.10 47.14 63.81 13.14 50.38 79.52
GaLR (Yuan et al. 2022b) 6.59 19.85 31.04 4.69 19.48 32.13 14.82 31.64 42.48 11.15 36.68 51.68
RSICD, RSITMD, and/or UCM-caption not seen in training
CLIP-original 6.59 20.68 31.75 3.62 14.28 23.63 10.18 30.31 42.04 8.31 24.96 39.03 37.62 78.10 89.52 28.12 64.99 77.19
CLIP-laion-RS 8.42 23.70 35.86 5.81 19.49 30.25 13.72 32.08 44.91 10.57 31.48 46.96 39.52 79.52 90.00 29.71 62.60 80.37
SkyCLIP-30 8.97 24.15 37.97 5.85 20.53 33.53 11.73 33.19 47.35 10.19 32.47 49.08 38.57 84.29 93.81 31.83 64.19 81.96
DOFA-CLIP-L 8.42 25.16 37.05 8.15 24.90 37.73 13.94 30.31 44.47 13.45 38.70 55.78 43.33 81.43 93.81 35.28 71.62 87.80
Overall zero-shot performance
As shown in Table 2, DOFA-CLIP establishes new state-of-the-art results in zero-shot remote-sensing image classification. Compared with existing CLIP-based baselines such as CLIP-original, RemoteCLIP, CLIP-laion-RS, and the various SkyCLIP configurations, DOFA-CLIP-B-224 consistently outperforms all counterparts. Its average accuracy across the eight scene datasets is significantly higher than that of the other base models. The Larger variant of DOFA-CLIP with higher resolution yields further accuracy gains. On three fine-grained roof-attribute classification tasks, both DOFA-CLIP variants achieve strong accuracy on roof shape and surface, significantly outperforming other methods. Although the results on smoothness remain more modest, our approach still provides a clear improvement over existing approaches.
Zero-shot classification on multispectral imagery
The zero-shot classification results on three GEO-Bench datasets are presented in Table 3. DOFA-CLIP demonstrates competitive or superior zero-shot multi-label classification performance compared to existing models on the m-bigearthnet dataset (43 classes). For the m-so2sat dataset (17 classes), when utilizing five spectral bands, the performance decreases slightly due to the significantly less multispectral training data compared to RGB. Note that competitive models cannot process 5-band images as input for zero-shot classification. On the m-forestnet dataset (12 classes), DOFA-CLIP substantially outperforms other models and shows remarkable accuracy improvements when using five spectral bands as input. Notably, the m-forestnet dataset uses Landsat-8 data, which is unseen in our training dataset.
Cross-modal retrieval performance
Table 4 compares DOFA-CLIP against recent methods on three benchmark caption datasets: RSICD, RSITMD, and UCM-caption. We report Recall@1, Recall@5, and Recall@10 for both image-to-text and text-to-image retrieval. Despite the inherent difficulty of remote-sensing cross-modal tasks, DOFA-CLIP-L consistently achieves competitive or higher recall scores than most competitive baselines, including CLIP-original, CLIP-laion-RS, and SkyCLIP-30.
Table 5: Ablation experiments of models on the AID zero-shot classification dataset.
Model Distill. Top-1 Accuracy Top-5 Accuracy
SigLIP 65.95 96.20
DOFA-CLIP-B (RGB) 76.90 94.55
SigLIP+DOFA-CLIP-B (RGB) 77.70 95.50
SigLIP+DOFA-CLIP-B (RGB) 72.00 95.80
DOFA-CLIP-B (Others) 64.70 91.95
SigLIP+DOFA-CLIP-B (Others) 69.50 96.20
SigLIP+DOFA-CLIP-B (Others) 66.30 91.20
DOFA-CLIP-B (Mixed) (baseline) 71.60 93.65
DOFA-CLIP-B (Merging) (ours) 77.60 96.65
Ablation studies
Table 6: Ablation experiments of different distillation methods on the segmentation datasets.
Model MADOS m-nz-cattle m-NeonTree
SigLIP 57.2 77.5 52.3
DOFA-CLIP-B w/o MaKA 61.7 81.5 58.1
DOFA-CLIP-B w/ MaKA 62.3 82.2 59.0
Table 7: Ratio for merging SigLIP and DOFA-CLIP-B (RGB).
$m_{1}$ AID (Top-1) RSI-CB (Top-1) Roof shape (Top-1)
0 65.95 38.01 41.17
0.1 44.25 27.33 37.67
0.3 56.75 31.87 38.00
0.5 74.65 41.80 44.33
0.7 77.40 45.36 53.67
0.9 77.70 47.03 57.83
1.0 76.90 46.72 58.50
Weight merging and MaKA
Table 5 presents the results of our ablation experiments, evaluating the effects of weight merging and MaKA across different models on the AID zero-shot classification dataset. The baseline SigLIP model can be significantly improved by incorporating the DOFA-CLIP-B (RGB) model, trained on the RGB part of GeoLangBind-2M. The model trained with our MaKA module (for distillation) can largely outperform the non-distilled version. DOFA-CLIP-B (Others), trained on non-RGB modalities, performs lower than its RGB counterpart. When merged with SigLIP, the performance can be significantly improved. DOFA-CLIP-B (Mixed) is the model trained with all the data modalities mixed together. When compared with the DOFA-CLIP-B (Merging), the performance is significantly lower. This demonstrates that weight merging is more effective and flexible than direct mixing of data modalities.
Figure 4: t-SNE visualization of feature distributions for SigLIP features (left) and DOFA-CLIP-L-384 model (right).
MaKA enhances segmentation performance
As presented in Table 6, we conducted ablation studies on the MADOS (Kikaki et al. 2024), m-nz-cattle (Abuaiadah and Switzer 2022), and m-NeonTree (Weinstein et al. 2021) datasets. DOFA-CLIP-B without MaKA represents a model trained using a baseline distillation method that does not incorporate modality-specific wavelengths as conditioning. The results demonstrate the significant effectiveness of the MaKA module, showing consistent performance improvements across all datasets. More details are in the supplementary material.
Weight merging ratios
In Table 7 and 8, a linear search is conducted to determine the optimal ratios for weight merging. Based on the results, we select 0.9 for $m_{1}$ to merge the weights of SigLIP and DOFA-CLIP (RGB). We choose 0.5 for merging weights of the intermediate weights and DOFA-CLIP (Others). Thus, we set $m_{1}=0.9$ and $m_{2}=0.5$ to derive our final model weights. Note that the weight merging is executed a single time using a designated zero-shot classification dataset, and held fixed across all tasks.
Table 8: Ratio for merging weights of DOFA-CLIP-B (RGB) and DOFA-CLIP-B (Others).
$m_{2}$ AID (Top-1) RSI-CB (Top-1) Roof shape (Top-1)
0 77.70 47.03 57.83
0.1 78.10 47.31 57.50
0.3 78.55 47.91 55.83
0.5 78.60 49.17 52.67
0.7 76.45 49.92 52.00
0.9 76.35 43.98 63.33
1.0 69.20 34.80 59.67
Figure 5: Visual comparison of deep features from RemoteCLIP, Skyscript, and DOFA-CLIP.
Feature visualization and analysis
On the left side of Fig. 4, we present the t-SNE plot of SigLIP features. Compared to the SigLIP features, the features of DOFA-CLIP-L-384 demonstrate a significantly stronger imagery-text alignment across all modalities, making it a powerful foundation for Earth observation tasks. In Fig. 5, we visually compare the feature heatmaps of RemoteCLIP, SkyScript, and DOFA-CLIP. DOFA-CLIP features are more spatially structured and preserve more object details, demonstrating the effectiveness of VECT.
Conclusion
We introduce DOFA-CLIP, a unified vision–language foundation model that bridges heterogeneous EO modalities through a shared language space. To achieve this, we construct GeoLangBind-2M, a large-scale image–text dataset containing six EO data modalities. DOFA-CLIP consists of a wavelength-aware encoder and a modality-aware knowledge agglomeration module to enhance fine-grained image understanding. Additionally, progressive weight merging is proposed to scale the model training to multiple data modalities. Extensive experiments demonstrate the state-of-the-art performance of DOFA-CLIP across zero-shot classification, semantic segmentation, and retrieval tasks.
Supplementary material
The supplementary material contains the following sections:
  1. 1.
    Details of the evaluation datasets;
  2. 2.
    Details on the construction of GeoLangBind-2M;
  3. 3.
    More visualization examples of GeoLangBind-2M;
  4. 4.
    Ablation studies on the distillation loss terms;
  5. 5.
    Confusion matrices of the zero-shot classification tasks;
  6. 6.
    More visualization of the features;
Details of the evaluation datasets
We evaluate DOFA-CLIP on both datasets that have been established in previous works for direct comparison, as well as additional datasets that highlight the wide-ranging capabilities of the proposed model.
The following datasets are part of the SkyScript (Wang et al. 2024c) evaluation framework, which we follow for direct comparison.
  • SkyScript classification dataset (Wang et al. 2024c): contains 7,000 aerial RGB images with 70 classes and different objects than the SkyScript pretraining dataset.
  • AID (Xia et al. 2017): contains 2,000 aerial RGB images and 30 image scene classes.
  • EuroSAT (Helber et al. 2019): contains 27,000 Sentinel-2 images with 13 spectral bands and 10 image scene classes. We use the dedicated test split that contains 2,700 images.
  • fMoW (Christie et al. 2018): contains 106,081 RGB aerial images with 62 image scene classes.
  • Million-AID (Long et al. 2021): contains 10,000 RGB aerial images of varying GSD with 51 image scene classes.
  • PatternNet (Zhou et al. 2018): contains 30,400 high-resolution RGB aerial images with 6–50 cm GSD and 38 image scene classes.
  • NWPU-RESISC45 (Cheng, Han, and Lu 2017): contains 31,500 aerial RGB images with 0.2–30 m GSD and 45 image scene classes.
  • RSI-CB (Li et al. 2017): contains 24,747 aerial RGB images and 35 image scene classes.
Table 9: GEO-Bench dataset description.
Tasks Name Image Size # Classes Train Val Test # Bands Sensors
Classification m-bigearthnet 120 x 120 43 20000 1000 1000 12 Sentinel-2
m-so2sat 32 x 32 17 19992 986 986 18 Sentinel-2 + Sentinel-1
m-forestnet 332 x 332 12 6464 989 993 6 Landsat-8
Segmentation m-nz-cattle 500 x 500 2 524 66 65 3 RGB
m-NeonTree 400 x 400 2 270 94 93 5 RGB + Hyperspectral + Elevation
Additionally, we choose the GEO-Bench (Lacoste et al. 2024) suite of datasets that cover a range of relevant remote sensing tasks across different domains, sensors, and geospatial locations. Table 9 contains a detailed overview of the number of samples, sensors, and target classes in the GEO-Bench datasets.
Figure 6: Word cloud for the Flair2-RGB-caption dataset.
Figure 7: Word histogram for the Flair2-RGB-caption dataset.
In addition, we conduct ablation experiments on the Marine Debris and Oil Spill (MADOS) (Kikaki et al. 2024) dataset. The MADOS dataset is a high-resolution multispectral Sentinel-2 dataset for marine pollution detection. Spanning 174 globally distributed scenes (2015–2022) with 1.5M annotated pixels, it covers diverse pollutants, sea surface features, and water-related classes. Unlike existing datasets, MADOS enables scalable, generalizable deep learning models for holistic marine pollution monitoring.
Details on constructing GeoLangBind-2M
We construct our multimodal image–text dataset GeoLangBind-2M by assembling existing datasets and curating 8 new datasets. Details on the newly curated sub-datasets are as follows.
Flair2-RGB-caption
For the Flair2-RGB-caption dataset, originally designed for semantic segmentation land cover classification, we employed an approach similar to the ChatEarthNet (Yuan et al. 2024) dataset construction. We select the training set of 61,711 images from FLAIR #2 (Garioud et al. 2023) and derive contextual information from their corresponding pixel-level segmentation masks. This extracted semantic content, paired with the segmentation maps as visual context, serves as the input for generating comprehensive descriptions using the Pixtral 12B (Agrawal et al. 2024) language model.
Our semantic extraction process involves calculating the distribution percentages of various terrain categories within each image and formulating these statistics into structured contextual cues. We identify all land cover categories in each segmentation map, assigning distinctive color codes to represent different semantic classes. The designed prompt explicitly defines the correspondence between color codes and land cover categories, while also incorporating the calculated percentage coverage of each land/object type within the segmentation maps. This quantitative spatial information is then processed by the Pixtral 12B model to analyze distribution patterns and generate linguistically accurate descriptions.
With this semantic framework established, we craft a prompt for generating detailed image descriptions. The prompt template variables indicated in brackets represent dynamic content populated during the semantic processing phase. The exact prompt template used for Flair2-RGB-caption creation is shown below:
{mdframed}
[backgroundcolor=gray!5, linecolor=black] You are an AI visual assistant capable of describing a scene based on a segmentation map. The map represents different land cover types using specific colors. The legend is as follows:
- [$color\_i$] corresponds to [$label\_i$], occupying [$percent\_i$]of the image.
Do not mention colors, color coding, or technical details. Use the given land cover class names exclusively. Generate a brief yet natural description of the scene by extending the sentence: “The aerial image contains [$presented\_labels$] land types.” Provide a concise summary of their spatial distribution.
In Fig. 6, we illustrate the word cloud for this dataset.
Flair2-Elevation-caption
The Flair2-Elevation-Caption dataset is constructed to generate detailed textual descriptions of terrain characteristics by leveraging both semantic segmentation maps and elevation maps. The dataset creation follows a structured approach, similar to the Flair2-RGB-caption dataset, ensuring that the generated captions effectively describe the topographical and land cover features of each image.
Figure 8: Word cloud for the Flair2-Elevation-caption dataset.
Figure 9: Histogram for the Flair2-elevation-caption dataset. Similarly to the other datasets of this appendix, FLAIR2-Elevation-caption has a focused vocabulary with many rare domain-specific terms. Its distribution deviates from a true power law due to frequent word repetition and an unusually long tail that is populated with remote sensing imagery terms. This mix of common and rare words could help models generalize by balancing domain-specific terms with diverse scene descriptors.
To begin, we select 61,711 images from the FLAIR #2 dataset, each accompanied by a segmentation map and an elevation map. The segmentation map provides land cover information, where each pixel corresponds to a specific land class, while the elevation map encodes height values representing terrain variation. Additionally, a class label mapping is used to associate numerical class IDs with their corresponding semantic labels, such as “forest,” “urban area,” or “water body.” This combination of datasets enables a comprehensive understanding of both land cover distribution and elevation patterns.
For each selected image, we analyze the elevation distribution by identifying the highest and lowest elevation values within the image patch. The corresponding land cover types at these extreme points are then determined using the segmentation map. This extraction process helps characterize the relationship between land cover and elevation, capturing how different landscapes correspond to varying altitude levels.
Using the extracted elevation and land cover information, we design a structured prompt to generate descriptive and natural textual captions. The prompt explicitly states the highest and lowest elevation values and their associated land cover types, followed by a request to summarize the overall terrain characteristics. The model is guided to describe whether the landscape is relatively flat or exhibits significant elevation changes, ensuring that the generated captions provide useful insights for downstream geospatial tasks.
To generate the final dataset, the structured prompts are input into a language model Pixtral 12B, which produces detailed, human-like descriptions of the elevation characteristics. These generated captions are then stored alongside their corresponding elevation and segmentation maps, forming the Flair2-Elevation-Caption dataset. The dataset offers rich, multimodal learning signals, allowing models to develop a deeper understanding of both land cover semantics and terrain variations.
{mdframed}
[backgroundcolor=gray!5, linecolor=black] This is an elevation map that indicates the height of each pixel. The highest areas, at an elevation of approximately [${highest\_height}$] meters, are [$highest\_class$]. The lowest areas, at an elevation of approximately [${lowest\_height}$] meters, are [$lowest\_class$].
Based on the provided context and elevation values, generate a concise and accurate description of the elevation map. Describe the image by briefly addressing:
1) The highest and lowest land cover types.
2) Whether the terrain is relatively flat or has significant elevation differences. In Fig. 8, we demonstrate the word cloud to provide a better overview of the dataset. The histogram of the Flair2-Elevation-caption dataset is presented in Fig. 9.
MMflood-SAR-caption
The MMflood-SAR-caption dataset contains textual descriptions of flood maps generated using Pixtral 12B. In the following, we will describe the generation process in detail. Each flood map is represented as a binary mask, where white pixels indicate flooded areas, and black pixels represent non-flooded regions. To create meaningful textual descriptions, we analyze the extent of flooding and its spatial distribution before constructing structured prompts to guide the language model.
Figure 10: Word cloud for the MMflood-SAR-caption dataset.
Figure 11: Word histogram for the MMflood-SAR-caption dataset.
To quantify the severity of flooding, we calculate the total number of flooded pixels in each binary mask and express it as a percentage of the total image area. This value provides a direct measure of the flood extent, allowing the generated captions to capture variations in flood severity across different samples. In addition to flood coverage, we assess spatial patterns by dividing the image into four quadrants: top, bottom, left, and right. By checking the presence of flooded pixels in each region, we determine whether the flooding is concentrated in specific areas or dispersed across the map.
Once these characteristics are extracted, we construct a structured prompt to guide Pixtral 12B in generating concise and informative descriptions. The prompt instructs the model to summarize the proportion of flooded areas and specify their locations within the image. To ensure clarity and brevity, the generated description is limited to fewer than 70 words, making it suitable for vision–language pretraining.
The prompts are then fed into Pixtral 12B, which generates natural language captions describing each flood map. These captions, stored alongside their corresponding binary masks, form the MMFlood dataset, enabling automated flood analysis, disaster monitoring, and geospatial AI research. By combining semantic flood mapping with natural language descriptions, this dataset enhances the ability of AI models to interpret and describe flood-affected regions in a human-readable format. We also provide the word cloud in Fig. 10 to show the detailed content of this dataset. In Fig. 11, we present the histogram of the words in the MMflood-SAR-caption dataset.
{mdframed}
[backgroundcolor=gray!5, linecolor=black] This is a flood map where areas marked with white pixels indicate flooded regions. The flooded area occupies approximately [$flood\_percentage$]of the entire map. Please analyze the flood map and provide insights into the affected areas.
You must generate a short description (less than 70 words) of the elevation image. First, describe the portion of floods; then introduce the location of the flooded areas.
After generating these detailed textual descriptions, we further refine the dataset by prompting Pixtral 12B to summarize each description into a concise 30-word caption. The prompts used in this summarizing process are as follows:
{mdframed}
[backgroundcolor=gray!5, linecolor=black] You are an AI assistant tasked with creating a concise 30-word caption that effectively summarizes the key points of the following content: [$caption$]
NLCD-hyper-caption
The caption generation process for hyperspectral images follows a methodology similar to the FLAIR #2 dataset, ensuring consistency in semantic information extraction and description generation. Hyperspectral images provide rich spectral information across multiple wavelengths, allowing for a more detailed classification of land cover types. However, the caption generation process remains aligned with the Flair2-RGB-Caption dataset, differing only in the specific prompt format used to guide the Pixtral 12B model.
Figure 12: Word cloud for the NLCD-hyper-caption dataset.
Figure 13: Word histogram for the NLCD-hyper-caption dataset.
To create captions, we first extract semantic information from the hyperspectral images using their corresponding segmentation maps. Each land cover type is identified and labeled, and its spatial distribution within the image is analyzed. Unlike conventional RGB imagery, where color-based segmentation might play a role, hyperspectral images rely on spectral signatures to differentiate land cover types, making this process crucial for accurate caption generation.
Once the land cover labels and their distributions are determined, we construct a structured prompt to generate natural language descriptions. The Pixtral 12B model is tasked with producing concise and human-readable descriptions of the scene. The prompt ensures that the generated captions focus solely on land cover types and their spatial distributions, avoiding any mention of technical details, color coding, or hyperspectral data intricacies.
Following the generation of detailed textual descriptions, we further prompt Pixtral 12B to refine the descriptions into brief and natural captions. This step ensures that the dataset provides both long-form and short-form textual annotations, improving its usability for vision–language tasks, remote sensing analysis, and multimodal learning.
The prompt template used for hyperspectral image captioning is as follows:
{mdframed}
[backgroundcolor=gray!5, linecolor=black] You are an AI visual assistant capable of describing a scene based on a segmentation map. The map represents different land cover types using specific colors. The legend is as follows:
- [$color\_i$] corresponds to [$label\_i$], occupying [$percent\_i$]of the image.
Do not mention any colors, color coding, or technical details. Use the given class names. Only mention land cover types in the color legend.
Generate a brief and natural description of the scene by refining: “The hyperspectral image contains [$presented\_labels$] land types.” Provide a concise description of their spatial distributions (e.g., left, right, top, bottom).
To provide a clearer overview of this dataset, we showcase the word cloud in Fig. 12. We also showcase the histogram of the words in the NLCD-hyper-caption dataset in Fig. 13.
SAR-ship-caption and IR-ship-caption
For the SAR-ship-caption and IR-ship-caption subsets, we derive short captions from existing datasets to ensure high-quality textual descriptions. For IR-ship-caption, we utilize samples from the HIT-UAV dataset, specifically the data provided in Zhang et al. (2024a), which is originally designed for visual question answering tasks. To generate descriptions for infrared images, we concatenate the corresponding question and answer pairs, transforming them into concise yet informative captions.
Similarly, for the SAR-ship-caption dataset, we generate captions for SAR images using three key datasets: HRSID (Wei et al. 2020), SSDD (Zhang et al. 2021), and AIR-SARShip-2.0 (Sun, Sun, and Wang 2020). These datasets provide extensive annotated SAR imagery, enabling the creation of meaningful captions. For further details on the captioning methodology, please refer to Zhang et al. (2024a).
ChatEarthNet captions
Figure 14: Word cloud for the ChatEarthNet-caption datasets.
Figure 15: Word histogram for the ChatEarthNet-caption datasets.
Regarding the ChatEarthNet-S2-caption and ChatEarthNet-SAR-caption datasets, we reuse the detailed descriptions provided in the original ChatEarthNet dataset. Specifically, we use Pixtral 12B (Agrawal et al. 2024) to summarize these detailed descriptions into short image captions and generate question-answer pairs for visual question answering tasks. The word cloud for ChatEarthNet-caption datasets is presented in Fig. 14. To provide more insights into the dataset, we also show the histogram in Fig. 15. The prompt is as follows: {mdframed}[backgroundcolor=gray!5, linecolor=black] You are an AI assistant tasked with creating a concise 30-word caption that effectively summarizes the key points of the following content: [$caption$]
More dataset examples
These paired samples not only capture scene-level descriptions but also highlight fine-grained properties, such as object positions and fine-grained land cover types. Thereby the dataset enables the development of foundation models proficient in downstream performance for land cover mapping, flood monitoring, and object detection under challenging conditions (e.g., night or cloudy weather).
In Fig. 16 and Fig. 17, we provide more examples to provide a better overview of our dataset.
Figure 16: More examples of the image-text pair of GeoLangBind-2M, our multimodal image-text dataset.
Figure 17: More examples of the image-text pair of GeoLangBind-2M, our multimodal image-text dataset.
Ablation studies on distillation loss terms
Table 10: Distillation loss balancing weights vs. accuracy
$\alpha_{s}$ (SigLIP) $\alpha_{d}$ (DinoV2) $\alpha_{v}$ (ViT) aid (Top-1) eurosat (Top-1)
0 0 0 59.10 32.96
1.0 0.0 0.0 61.42 33.63
1.0 1.0 0.0 65.36 35.26
1.0 1.0 1.0 68.75 38.63
2.0 0.0 0.0 64.85 33.19
2.0 0.5 0.5 64.70 25.30
2.0 1.0 1.0 69.20 36.67
Table 10 presents an ablation study on the impact of different distillation loss balancing weights ($\alpha_{s}$, $\alpha_{d}$, $\alpha_{v}$) on classification accuracy for the AID and EuroSAT datasets. The baseline model (no distillation) achieves 59.10% (AID) and 32.96% (EuroSAT). Adding SigLIP distillation ($\alpha_{s}=1.0$) improves performance, while combining SigLIP and DinoV2 ($\alpha_{s}=1.0,\alpha_{d}=1.0$) further enhances results, especially for AID (65.36%). Introducing ViT distillation ($\alpha_{v}=1.0$) boosts EuroSAT accuracy, peaking at 38.63% when $\alpha_{s}=\alpha_{d}=\alpha_{v}=1.0$. In this work, we adopt the best overall configuration, $\alpha_{s}=2.0,\alpha_{d}=1.0,\alpha_{v}=1.0$, achieves the highest AID accuracy (69.20%) and strong EuroSAT performance (36.67%). This demonstrates the effectiveness of multi-teacher distillation.
Confusion matrices of zero-shot classification tasks
Figure 18 to Fig. 25 present the confusion matrices of our DOFA-CLIP-L-384 model on all eight datasets for zero-shot classification. These confusion matrices reveal that common model mistakes largely occur for classes that humans would also have trouble distinguishing, or for classes with similar semantic meanings. For instance, the model commonly confuses “sparse residential” and “medium residential” (RESISC45), “annual crop” and “permanent crop” (EuroSAT), “aquaculture land” and “pond” (SkyScript), “chaparral” and “desert” (RESISC45), “mine” and “quarry” (SkyScript, Million-AID), “recreational facility” and “stadium” and “playground” (fMoW, AID), and “ship” and “harbor” and “ferry terminal” (RESISC45, PatternNet).
In Fig. 26, we show the comparison of confusion matrices between SkyScript (ViT-L) and DOFA-CLIP-L-384 on fine-grained zero-shot classification tasks. It can be clearly seen that our model performs better.
Figure 18: Confusion matrix of the zero-shot classification results of DOFA-CLIP-L-384 on SkyScript dataset.
Figure 19: Confusion matrix of the zero-shot classification results of DOFA-CLIP-L-384 on AID dataset.
Figure 20: Confusion matrix of the zero-shot classification results of DOFA-CLIP-L-384 on EuroSAT dataset.
Figure 21: Confusion matrix of the zero-shot classification results of DOFA-CLIP-L-384 on fMoW dataset.
Figure 22: Confusion matrix of the zero-shot classification results of DOFA-CLIP-L-384 on Million-AID dataset.
Figure 23: Confusion matrix of the zero-shot classification results of DOFA-CLIP-L-384 on PatternNet dataset.
Figure 24: Confusion matrix of the zero-shot classification results of DOFA-CLIP-L-384 on the RESISC45 dataset.
Figure 25: Confusion matrix of the zero-shot classification results of DOFA-CLIP-L-384 on RSICB dataset.
Figure 26: Comparison of confusion matrices between SkyScript-Large and DOFA-CLIP-L-384 on the fine-grained zero-shot classification task.
More visualization of features
To better understand the representations learned by different models, Fig. 27 visualizes the feature maps extracted from RemoteCLIP (ViT-L), SkyScript (ViT-L), and DOFA-CLIP-L-384 across two datasets: m-chesapeake and m-nz-cattle from GEO-Bench. The m-chesapeake dataset primarily focuses on land cover classification, while the m-nz-cattle dataset captures cattle in remote sensing images. The feature visualization highlights key differences in spatial structure, consistency, and detail preservation across models.
Spatially structured features
DOFA-CLIP produces features with higher spatial resolution, enabling finer spatial distinctions compared to RemoteCLIP and SkyScript. The feature maps exhibit well-defined structures that align with meaningful spatial regions in the original images, such as roads, buildings, and vegetation.
Consistency across semantic classes
The feature maps from DOFA-CLIP demonstrate greater consistency across images with similar semantic content. As observed, buildings and roads consistently exhibit high activations, while tree-covered areas show lower activations. This structured activation pattern is apparent across multiple samples, suggesting that DOFA-CLIP learns more stable and semantically aligned representations.
Preservation of details
DOFA-CLIP retains finer details within its feature maps, capturing small-scale variations and object boundaries more effectively than the baselines. This advantage is particularly evident in the m-nz-cattle dataset, where cattle are localized with sharper feature activations, as highlighted by the red bounding boxes. RemoteCLIP and SkyScript, on the other hand, exhibit more diffused and less distinct feature responses, making object-level distinctions less precise.
Figure 27: Visualization of features from RemoteCLIP (ViT-L), SkyScript (ViT-L), and DOFA-CLIP-L-384 on the m-chesapeake and m-nz-cattle datasets from GEO-Bench.